# Make Every Penny Count: Difficulty-Adaptive Self-Consistency for Cost-Efficient Reasoning

### v2

@@ A.1 Prompt for Difficulty Ranking
| Method | Input tokens | Output tokens | Cost | Acc |
| SC | 846 | 6045 | 0.3881 | 73.19 |
| ASC | 15982 | 3265 | 0.6754 | 73.14 |
| ESC | 4458 | 4415 | 0.3986 | 73.18 |
| DSC | 1617 | 3857 | 0.2799 | 73.18 |
A.2 Generalizability of DSC On Small Open-source Language Model
To further explore the effectiveness of DSC on small open-source models, we conduct experiments on the open-source model Mistral-7B-Instruct-v0.3 (Jiang et al., 2023) using GSM8K dataset. As shown in Table 8, we convert the token cost into price according to that of GPT-4 for a simpler comparison and keep the other settings completely consistent with the main experiment (Please note that for DSC, the cost of step 1 and step 2 is included.). The experimental results indicate that DSC has the potential to work effectively on smaller models.
A.3 Comparison of Inference Time Between Different Methods
@@
A.4 Comparison with ESC under Different Window Size
@@
A.5 Hyperparameter Experiment of Difficulty Ranking
@@
A.6 Background of ESC and ASC
@@
ASC proposes the Dirichlet Stopping Criteria, where sampling is conducted one by one. After each sampling, the Dirichlet Stopping Criteria is used to determine whether all current samples meet a specific distribution. If they do, the sampling is stopped. If not, the one-by-one sampling continues until the preset maximum sampling value is reached. The Dirichlet Stopping Criteria is shown in Equation 1, where represents the current set of samples, is the number of , is the preset threshold (set to 0.95 for ASC in default), and is the probability of the most frequently occurring answer in the set . The core idea of ASC is to measure the two answers with the highest probability in the current set of samples. If the difference between the probability of the answer with the highest probability and the second highest probability exceeds a threshold (), it can be considered that the model is very confident in the answer with the highest probability, and sampling can be stopped.

### v3

@@ A.1 Prompt for Difficulty Ranking
A.2 Comparison of Inference Time Between Different Methods
@@
A.3 Comparison with ESC under Different Window Size
@@
A.4 Performance of Difficulty Ranking under Mixing Types of Questions
To explore whether Difficulty Ranking can effectively rank problems of different types in terms of difficulty, we compare the performance of Difficulty Ranking under mixed and unmixed problem types on MATH, as shown in Table 10. Specifically, for unmixed, we perform Difficulty Ranking on seven subsets separately; for mixed, we directly execute Difficulty Ranking on random shuffled MATH dataset, and calculate the correlation on the entire dataset for both. The results indicate that mixing different types of questions for LLM to rank can lead to a decrease in performance. This might be due to the challenge for LLM in adhering to the same assessment standards for various types of questions.101010It could also lead to inaccurate ranking for humans.
@@ A.4 Performance of Difficulty Ranking under Mixing Types of Questions
| Model | Metric | Unmixed | Mixed |
| GPT-4 | Spearman | 0.6451 | 0.5148 |
| Pearson | 0.6487 | 0.5197 | |
| Kendall | 0.5060 | 0.3931 | |
| GPT-3.5-Turbo | Spearman | 0.4828 | 0.4070 |
| Pearson | 0.4887 | 0.4145 | |
| Kendall | 0.3697 | 0.3073 |
| Setting | N | ||||
| 1 | 5 | 10 | 20 | 40 | |
| Zero-shot | 27.01 | 31.44 | 33.84 | 35.21 | 35.93 |
| Few-shot | 33.08 | 38.54 | 41.39 | 43.01 | 43.94 |
A.5 Hyperparameter Experiment of Difficulty Ranking
A.6 Necessity of Few-Shot Setting in DSC
Previous studies (Wei et al., 2022a; Wang et al., 2023) have experimentally demonstrated that compared to zero-shot (without demonstrations), few-shot (with demonstrations) significantly improves the performance of LLMs on reasoning tasks. Therefore, we adopted the same few-shot setting as SC, ASC, and ESC. To further verify this, we test the Accuracy of SC on GPT-3.5-Turbo under both zero-shot and few-shot (8-shot as default) settings with different sample sizes N on MATH500.
The results (Table 11) show that removing demonstrations leads to a significant performance drop, validating the necessity of multiple demonstrations. As a matter of fact, during our research, we tried to dynamically allocate the number of demonstrations based on question difficulty to reduce input costs. However, we find that reducing the number of demonstrations significantly decreased performance, regardless of the difficulty of the questions.
Moreover, recent study (Agarwal et al., 2024) has demonstrated that a greater number of demonstrations generally leads to more performance improvements. For most tasks, the optimal number of demonstrations is often greater than 512, far exceeding the default setting of 8 in DSC, which results in much higher input costs. This further highlights the value of DSC.
A.7 Comparison of DSC with Self-Correction and Verifier-Based Methods
We consider that self-correct (Wu et al., 2024b), verifier-based methods (Wu et al., 2024a), and Self-consistency (Wang et al., 2023) can all be seen as approaches that aim to improve model performance by increasing computational effort, albeit through different implementation paths. In comparison, DSC takes a different direction by focusing on optimizing efficiency—reducing computational costs as much as possible while maintaining model performance. This distinction positions DSC as complementary to self-correct, verifier-based methods, and Self-consistency, with a unique focus on efficiency.
Furthermore, while DSC is implemented based on Self-consistency, its concept of dynamically allocating computational resources according to question difficulty could also be adapted to self-correct and verifier-based methods to enhance efficiency. For example, in the case of self-correct, this might involve assessing whether a query requires self-correction based on its difficulty and estimating the necessary degree of correction (e.g., the number of iterations). Similarly, for verifier-based methods, which already involve multiple sampling followed by verification, integrating DSC could potentially lead to more efficient inference processes.
A.8 Background of ESC and ASC
@@
ASC (Aggarwal et al., 2023) proposes the Dirichlet Stopping Criteria, where sampling is conducted one by one. After each sampling, the Dirichlet Stopping Criteria is used to determine whether all current samples meet a specific distribution. If they do, the sampling is stopped. If not, the one-by-one sampling continues until the preset maximum sampling value is reached. The Dirichlet Stopping Criteria is shown in Equation 1, where represents the current set of samples, is the number of , is the preset threshold (set to 0.95 for ASC in default), and is the probability of the most frequently occurring answer in the set . The core idea of ASC is to measure the two answers with the highest probability in the current set of samples. If the difference between the probability of the answer with the highest probability and the second highest probability exceeds a threshold (), it can be considered that the model is very confident in the answer with the highest probability, and sampling can be stopped. For further details, please refer to the original paper.
A.9 Application Scenarios of DSC
As discussed in the Limitations section, DSC could be constrained in scenarios where a single user is restricted to one input at a time. However, it is widely applicable to common scenarios including: (1) servers (e.g., OpenAI or similar providers) managing a large number of simultaneous query requests from different users (e.g., ChatGPT receives requests from multiple users to solve multiple math problems simultaneously), and (2) cases where a user perform inference on a batch of data (E.g., Submitting multiple requests at once to efficiently receive feedback).
For these application scenarios, Recently, an increasing number of studies have focused on performing inference on large volumes of samples with LLMs (Cheng et al., 2023; Lin et al., 2024; Son et al., 2024; Liu et al., 2024a; Cong et al., 2025), which is computationally and financially costly in industrial and real-world applications. Therefore, the application scenarios of DSC have become a key focus in current research, attracting extensive attention from the academic and industrial communities, and we believe DSC’s potential applications and value will be more and more substantial in the future.